# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

### v1

@@ 4.2 Proofs of the Error Bound
where the third inequality follows from (56) and $\bm{W}_{l}^{*}=\bm{Q}_{l+1}\bm{\Sigma}_{l}^{*}\bm{Q}_{l}^{T}$ and $\bm{\Sigma}_{l}^{*}=\mathrm{BlkDiag}(\bm{\sigma}^{*},\bm{0})$ for each $\bm{Q}_{l}\in\mathcal{O}^{d_{l}-1}$ with $l=2,\dots,L$ according to 3. Similarly, we have $\|\bm{W}_{l}\|\leq 3\sigma^{*}_{\max}/2$ for each $l\in[L]$. Therefore, we have (57). Using Weyl’s inequality (see [17, Corollary 7.3.5]), we have for each $l\in[L]$ and $i\in[r_{\sigma}]$,
$\left|\sigma_{i}(\bm{W}_{l})-\sigma_{i}^{*}\right|\leq\|\bm{W}_{l}-\bm{W}_{l}^{*}\|\leq\frac{\sigma_{\min}^{*}}{2},$
which implies
$\sigma_{i}(\bm{W}_{l})\geq\sigma_{i}^{*}-\frac{\sigma_{\min}^{*}}{2}\geq\frac{\sigma_{\min}^{*}}{2}.$
@@
$\frac{1}{2}\nabla_{\bm{W}_{l}}G(\bm{W})\bm{W}_{l}^{T}=\bm{W}_{L:l+1}^{T}\left(\bm{W}_{L:1}-\sqrt{\lambda}\bm{Y}\right)\bm{W}_{l:1}^{T}+\lambda\bm{W}_{l}\bm{W}_{l}^{T},$
$\frac{1}{2}\bm{W}_{l+1}^{T}\nabla_{\bm{W}_{l+1}}G(\bm{W})=\bm{W}_{L:l+1}^{T}\left(\bm{W}_{L:1}-\sqrt{\lambda}\bm{Y}\right)\bm{W}_{l:1}^{T}+\lambda\bm{W}_{l+1}^{T}\bm{W}_{l+1}.$
Then, we have for each $l=1,\dots,L-1$,
$\frac{1}{2}\left(\nabla_{\bm{W}_{l}}G(\bm{W})\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\nabla_{\bm{W}_{l+1}}G(\bm{W})\right)=\lambda\left(\bm{W}_{l}\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\bm{W}_{l+1}\right).$
@@
$\left\|\bm{W}_{l}\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\bm{W}_{l+1}\right\|_{F}$ $=\frac{1}{2\lambda}\left\|\nabla_{\bm{W}_{l}}G(\bm{W})\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\nabla_{\bm{W}_{l+1}}G(\bm{W})\right\|_{F}$
$\leq\frac{3\sigma_{\max}^{*}}{4\lambda}\left(\|\nabla_{\bm{W}_{l}}G(\bm{W})\|_{F}+\|\nabla_{\bm{W}_{l+1}}G(\bm{W})\|_{F}\right)\leq\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda}\|\nabla G(\bm{W})\|_{F},$
where the first inequality follows from the triangle inequality, $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|\|\bm{B}\|_{F}$ for all $\bm{A},\bm{B}$ of the same size, and (57).
(iii) Using Weyl’s inequality and (59), we obtain for each $l\in[L]$ and $i\in[r_{\sigma}]$,
$|\sigma_{i}^{2}(\bm{W}_{l})-\sigma_{i}^{2}(\bm{W}_{l+1})|\leq\left\|\bm{W}_{l}\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\bm{W}_{l+1}\right\|_{F}\leq\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda}\|\nabla G(\bm{W})\|_{F}.$
This implies
$\left|\sigma_{i}(\bm{W}_{l})-\sigma_{i}(\bm{W}_{l+1})\right|$ $\leq\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda\left(\sigma_{i}(\bm{W}_{l})+\sigma_{i}(\bm{W}_{l+1})\right)}\|\nabla G(\bm{W})\|_{F}\leq\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F},$
where the last inequality follows from (58).
@@
$\bm{W}_{j:i}^{T}\bm{W}_{j:i}-(\bm{W}^{T}_{i}\bm{W}_{i})^{j-i+1}$ $=\bm{W}_{j-1:i}^{T}\left(\bm{W}_{j}^{T}\bm{W}_{j}-\bm{W}_{j-1}\bm{W}_{j-1}^{T}\right)\bm{W}_{j-1:i}$
$\quad+\bm{W}_{j-2:i}^{T}\left((\bm{W}_{j-1}^{T}\bm{W}_{j-1})^{2}-(\bm{W}_{j-2}\bm{W}_{j-2}^{T})^{2}\right)\bm{W}_{j-2:i}+\cdots$
$\quad+\bm{W}_{i}^{T}\left((\bm{W}_{i+1}^{T}\bm{W}_{i+1})^{j-i}-\bm{W}_{i}\bm{W}_{i}^{T})^{j-i}\right)\bm{W}_{i}$
$=\sum_{k=1}^{j-i}\bm{W}_{j-k:i}^{T}\left((\bm{W}_{j-k+1}^{T}\bm{W}_{j-k+1})^{k}-(\bm{W}_{j-k}\bm{W}_{j-k}^{T})^{k}\right)\bm{W}_{j-k:i}.$ (120)
@@
where the last inequality uses (59) and the second inequality follows from
=\ = $\left\|\sum_{l=1}^{k}(\bm{W}_{j+1-k}^{T}\bm{W}_{j+1-k})^{k-l}(\bm{W}_{j+1-k}^{T}\bm{W}_{j+1-k}-\bm{W}_{j-k}\bm{W}_{j-k}^{T})(\bm{W}_{j-k}\bm{W}_{j-k}^{T})^{l-1}\right\|_{F}$
[…]
This, together with (4.2), yields
@@
Using the same argument, we prove (62). Then, we complete the proof. ∎
Proof of 2.
@@
$\left\|(\bm{W}_{l-1}\bm{W}_{l-1}^{T})^{l-1}-(\bm{W}_{l}^{T}\bm{W}_{l})^{l-1}\right\|_{F}\leq\left\|(\bm{W}_{l-1}\bm{W}_{l-1}^{T})^{l-2}(\bm{W}_{l-1}\bm{W}_{l-1}^{T}-\bm{W}_{l}^{T}\bm{W}_{l})\right\|_{F}+\cdots$
$+\left\|(\bm{W}_{l-1}\bm{W}_{l-1}^{T}-\bm{W}_{l}^{T}\bm{W}_{l})(\bm{W}_{l}^{T}\bm{W}_{l})^{l-2}\right\|_{F}\leq\frac{3\sqrt{2}(l-1)\sigma_{\max}^{*}}{4\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2l-4}\|\nabla G(\bm{W})\|_{F},$
where the last inequality uses (57) and (59) in Lemma 3. This, together with the triangular inequality and (62), yields for $l=2,3,\dots,L$,
@@
Using the same argument, we have for $l=1,2,\dots,L-1$,
@@
For each $l\in[L]$, we compute
$\frac{1}{2}\nabla_{\bm{W}_{l}}G(\bm{W})$ =(13)${=}\bm{W}_{L:l+1}^{T}(\bm{W}_{L}\cdots\bm{W}_{1}-\sqrt{\lambda}\bm{Y})\bm{W}_{l-1:1}^{T}+\lambda\bm{W}_{l}=(\bm{W}_{l}\bm{W}^{T}_{l})^{L-1}\bm{W}_{l}-$
$\sqrt{\lambda}\bm{W}_{L:l+1}^{T}\bm{Y}\bm{W}_{l-1:1}^{T}+\lambda\bm{W}_{l}+\bm{R}_{2}(l)\bm{W}_{l}\bm{W}_{l-1:1}\bm{W}_{l-1:1}^{T}+\left(\bm{W}_{l}\bm{W}_{l}^{T}\right)^{L-l}\bm{W}_{l}\bm{R}_{1}(l).$
This, together with the triangle inequality, (4.2), and (122), yields for each $l\in[L]$,
@@
which directly implies (63). ∎
Next, using Lemma 3 and the SVD of $\bm{W}_{l}$ in (66), we proceed to the proof of 4.
Proof of 4.
Let $\bm{W}^{*}=(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}_{\bm{\sigma}^{*}}$ be such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})=\|\bm{W}-\bm{W}^{*}\|_{F}$. For ease of exposition, let $\bm{H}_{l}:=\bm{U}_{l-1}^{T}\bm{V}_{l}$ for each $l\in\{2,\dots,L\}$. According to (66), the $(i,j)$-th block of $\bm{H}_{l}$, denoted by $\bm{H}_{l}^{(i,j)}$, is
According to (67) with $\delta_{\sigma}\leq\sigma_{\min}^{*}$ and Lemma 3, we obtain (59). Substituting (66) and $\bm{H}_{l}=\bm{U}_{l-1}^{T}\bm{V}_{l}$ into (59) yields
$\left\|\bm{H}_{l}\bm{\Sigma}_{l}^{T}\bm{\Sigma}_{l}-\bm{\Sigma}_{l-1}\bm{\Sigma}_{l-1}^{T}\bm{H}_{l}\right\|_{F}\leq\frac{3\sqrt{2}\sigma^{*}_{\max}}{4\lambda}\|\nabla G(\bm{W})\|_{F}.$
Using this and the block structures of $\bm{\Sigma}_{l}$ in (66) and $\bm{H}_{l}$ in (124), we obtain
$\ \sum_{j=1}^{p+1}\sum_{i=1,i\neq j}^{p+1}\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}-\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\|_{F}^{2}$
$\leq$ $\ \sum_{j=1}^{p+1}\sum_{i=1}^{p+1}\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}-\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\|_{F}^{2}\leq\left(\frac{3\sqrt{2}\sigma^{*}_{\max}}{4\lambda}\right)^{2}\|\nabla G(\bm{W})\|_{F}^{2}.$ (125)
For each $l\in[L]$ and $i\in[p+1]$, we have $\|\bm{\Sigma}_{l,i}-\sigma^{*}_{t_{i}}\bm{I}\|\leq\|\bm{W}_{l}-\bm{W}_{l}^{*}\|\leq\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq{\delta_{\sigma}}/{3},$ where the first inequality follows from Weyl’s inequality and the last inequality uses (67). This implies
$\sigma^{*}_{t_{i}}-\frac{\delta_{\sigma}}{3}\leq\lambda_{\min}(\bm{\Sigma}_{l,i})\leq\lambda_{\max}(\bm{\Sigma}_{l,i})\leq\sigma^{*}_{t_{i}}+\frac{\delta_{\sigma}}{3}.$ (126)
For each $l\in[L]$ and $i,j\in[p+1]$ with $i>j$, we compute
$\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}\|_{F}\geq\lambda_{\min}^{2}(\bm{\Sigma}_{l}^{(j)})\|\bm{H}_{l}^{(i,j)}\|_{F}\geq\left(\sigma^{*}_{t_{j}}-\frac{\delta_{\sigma}}{3}\right)^{2}\|\bm{H}_{l}^{(i,j)}\|_{F},$ (127)
$\|\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\|_{F}^{2}\leq\lambda_{\max}\left(\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\right)\|\bm{H}_{l}^{(i,j)}\|_{F}\leq\left(\sigma^{*}_{t_{i}}+\frac{\delta_{\sigma}}{3}\right)^{2}\|\bm{H}_{l}^{(i,j)}\|_{F}.$ (128)
@@
$\quad\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}-\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\|_{F}\geq\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}\|_{F}-\|\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\|_{F}$
$\geq\left(\left(\sigma^{*}_{t_{j}}-\frac{\delta_{\sigma}}{3}\right)^{2}-\left(\sigma^{*}_{t_{i}}+\frac{\delta_{\sigma}}{3}\right)^{2}\right)\|\bm{H}_{l}^{(i,j)}\|_{F}\geq\frac{2\delta_{\sigma}\sigma^{*}_{\min}}{3}\|\bm{H}_{l}^{(i,j)}\|_{F},$
where the second equality follows from (127) and (128), and the last inequality is due to (48). Applying the same argument to the case where $i<j$, we obtain the same result. These, together with (4.2), yield
@@
$\|\bm{H}_{l}^{(i,i)}-\bm{P}_{l}^{(i)}\bm{Q}_{l}^{(i)^{T}}\|_{F}^{2}$ $=\sum_{k=1}^{g_{i}}\left(1-\sigma_{k}(\bm{H}_{l}^{(i,i)})\right)^{2}\leq\sum_{k=1}^{g_{i}}\left(1-\sigma_{k}^{2}(\bm{H}_{l}^{(i,i)})\right)^{2}$

### v3

@@ 3.2.1 Preliminary Results
where the third inequality follows from 1 and $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})<{\sigma^{*}_{\min}}/{2}$. Similarly, we have $\|\bm{W}_{l}\|\leq 3\sigma^{*}_{\max}/2$ for each $l\in[L]$. Using Weyl’s inequality (see, e.g., [18, Corollary 7.3.5]), we have $\left|\sigma_{i}(\bm{W}_{l})-\sigma_{i}^{*}\right|\leq\|\bm{W}_{l}-\bm{W}_{l}^{*}\|\leq\sigma_{\min}^{*}/{2}$ for all $i\in[r_{\sigma}]$ and $l\in[L]$. This implies $\sigma_{i}(\bm{W}_{l})\geq\sigma_{i}^{*}-{\sigma_{\min}^{*}}/{2}\geq{\sigma_{\min}^{*}}/{2}.$
@@
$\nabla_{l}G(\bm{W})\bm{W}_{l}^{T}=2\bm{W}_{L:l+1}^{T}\left(\bm{W}_{L:1}-\sqrt{\lambda}\bm{Y}\right)\bm{W}_{l:1}^{T}+2\lambda\bm{W}_{l}\bm{W}_{l}^{T},$
$\bm{W}_{l+1}^{T}\nabla_{l+1}G(\bm{W})=2\bm{W}_{L:l+1}^{T}\left(\bm{W}_{L:1}-\sqrt{\lambda}\bm{Y}\right)\bm{W}_{l:1}^{T}+2\lambda\bm{W}_{l+1}^{T}\bm{W}_{l+1}.$
Then, we have $\left(\nabla_{l}G(\bm{W})\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\nabla_{l+1}G(\bm{W})\right)/2=\lambda\left(\bm{W}_{l}\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\bm{W}_{l+1}\right)$ for each $l\in[L-1]$. This implies
$\left\|\bm{W}_{l}\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\bm{W}_{l+1}\right\|_{F}$ $=\frac{1}{2\lambda}\left\|\nabla_{l}G(\bm{W})\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\nabla_{l+1}G(\bm{W})\right\|_{F}$
$\leq\frac{3\sigma_{\max}^{*}}{4\lambda}\left(\|\nabla_{l}G(\bm{W})\|_{F}+\|\nabla_{l+1}G(\bm{W})\|_{F}\right)\leq\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda}\|\nabla G(\bm{W})\|_{F},$
where the first inequality follows from the triangle inequality, $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|\|\bm{B}\|_{F}$, and (35).
(iii) Using Weyl’s inequality and (36), we obtain that for each $l\in[L]$ and $i\in[r_{\sigma}]$,
$\left|\sigma_{i}^{2}(\bm{W}_{l})-\sigma_{i}^{2}(\bm{W}_{l+1})\right|\leq\left\|\bm{W}_{l}\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\bm{W}_{l+1}\right\|_{F}\leq\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda}\|\nabla G(\bm{W})\|_{F}.$
@@
[…]
@@
$\textstyle\bm{W}_{j:i}^{T}\bm{W}_{j:i}-(\bm{W}^{T}_{i}\bm{W}_{i})^{j-i+1}=\sum_{k=1}^{j-i}\bm{W}_{j-k:i}^{T}\left((\bm{W}_{j-k+1}^{T}\bm{W}_{j-k+1})^{k}-(\bm{W}_{j-k}\bm{W}_{j-k}^{T})^{k}\right)\bm{W}_{j-k:i}.$ (40)
@@
$\leq$ $\left\|(\bm{W}_{j+1-k}^{T}\bm{W}_{j+1-k})^{k}-(\bm{W}_{j-k}\bm{W}_{j-k}^{T})^{k}\right\|_{F}\prod_{l=i}^{j-k}\|\bm{W}_{l}\|^{2}$
$\leq$ $k\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2(j-i)}\left\|\bm{W}_{j+1-k}^{T}\bm{W}_{j+1-k}-\bm{W}_{j-k}\bm{W}_{j-k}^{T}\right\|_{F}$
[…]
$=\frac{\sqrt{2}k}{2\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2(j-i)+1}\|\nabla G(\bm{W})\|_{F},$
where the second inequality follows from
@@
This, together with (40), yields
@@
Using the same argument, we can prove (39). ∎
Corollary 2.
Under the setting of Lemma 2, we have
$\left\|(\bm{W}_{l}\bm{W}^{T}_{l})^{L-1}\bm{W}_{l}-\sqrt{\lambda}\bm{W}_{L:l+1}^{T}\bm{Y}\bm{W}_{l-1:1}^{T}+\lambda\bm{W}_{l}\right\|_{F}\leq c_{1}\|\nabla G(\bm{W})\|_{F},\ \forall l\in[L],$ (41)
where
Proof.
@@
$\left\|(\bm{W}_{l-1}\bm{W}_{l-1}^{T})^{l-1}-(\bm{W}_{l}^{T}\bm{W}_{l})^{l-1}\right\|_{F}$
$\leq\quad$ $\left\|(\bm{W}_{l-1}\bm{W}_{l-1}^{T})^{l-2}(\bm{W}_{l-1}\bm{W}_{l-1}^{T}-\bm{W}_{l}^{T}\bm{W}_{l})\right\|_{F}+\cdots+\left\|(\bm{W}_{l-1}\bm{W}_{l-1}^{T}-\bm{W}_{l}^{T}\bm{W}_{l})(\bm{W}_{l}^{T}\bm{W}_{l})^{l-2}\right\|_{F}$
[…]
This, together with the triangle inequality and (39), yields that for $l=2,3,\dots,L$,
@@
Using the same argument and (38), we have for $l=1,2,\dots,L-1$,
@@
For each $l=1,2,\ldots L$, we compute
[…]
$=$~ $(\bm{W}_{l}\bm{W}^{T}_{l})^{L-1}\bm{W}_{l}-\sqrt{\lambda}\bm{W}_{L:l+1}^{T}\bm{Y}\bm{W}_{l-1:1}^{T}$
$+\lambda\bm{W}_{l}+\bm{R}_{2}(l)\bm{W}_{l}\bm{W}_{l-1:1}\bm{W}_{l-1:1}^{T}+\left(\bm{W}_{l}\bm{W}_{l}^{T}\right)^{L-l}\bm{W}_{l}\bm{R}_{1}(l).$
This, together with the triangle inequality, (35), (3.2.1), and (44), yields for each $l\in[L]$,
@@
$\leq\$$\left(\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}\frac{(L-l)(L-l+1)+(l-1)l}{2\sqrt{2}\lambda}+\frac{1}{2}\right)\|\nabla G(\bm{W})\|_{F},$
which directly implies (41). ∎
To handle the repeated singular values in $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$, let $p\geq 1$ denote the number of distinct positive singular values. Recall from (34) that $r_{\sigma}$ denotes the number of positive singular values of $\bm{\sigma}^{*}$. Then there exist indices $t_{0},t_{1},\cdots,t_{p}$ such that $0=t_{0}<t_{1}<\cdots<t_{p}=r_{\sigma}$ and
Let $g_{i}:=t_{i}-t_{i-1}$ be the multiplicity of the $i$-th largest positive value of $\bm{\sigma}^{*}$ for each $i\in[p]$ and $g_{\max}:=\max\{g_{i}:i\in[p]\}$. Obviously, we have $r_{\sigma}=\sum_{i=1}^{p}g_{i}$ . Moreover, let
$\bm{W}_{l}$ $=\bm{U}_{l}\bm{\Sigma}_{l}\bm{V}_{l}^{T}$
$=\begin{bmatrix}\bm{U}_{l}^{(1)}&\!\!\!\!\dots\!\!\!&\bm{U}_{l}^{(p)}&\!\!\!\!\bm{U}_{l}^{(p+1)}\end{bmatrix}\mathrm{BlkD}\left(\bm{\Sigma}_{l}^{(1)},\dots,\bm{\Sigma}_{l}^{(p)},\bm{\Sigma}_{l}^{(p+1)}\right){\begin{bmatrix}\bm{V}_{l}^{(1)}&\!\!\!\!\!\!\dots\!\!\!\!\!&\!\bm{V}_{l}^{(p)}&\!\!\!\!\!\!\bm{V}_{l}^{(p+1)}\end{bmatrix}}^{T}$ (46)
With the above setup, we are ready to show that, for any point in the neighborhood of the set of critical points, the product of $\bm{U}_{l-1}$ and $\bm{V}_{l}$ is close to a block diagonal matrix with orthogonal diagonal blocks. Recall that $\delta_{\sigma}$ and $\mathcal{A}_{\rm sort}$ are defined in (28) and (29), respectively.
Proposition 3.
Suppose that Assumption 1 holds. Let $\bm{W}=(\bm{W}_{1},\dots,\bm{W}_{L})$ and $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary such that
$\left\|\bm{U}_{l-1}^{T}\bm{V}_{l}-\mathrm{BlkD}\left(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)}\right)\right\|_{F}\leq\frac{9\sigma_{\max}^{*}}{4\delta_{\sigma}\lambda\sigma^{*}_{\min}}\|\nabla G(\bm{W})\|_{F}.$ (48)
where $\bm{U}_{l}$ and $\bm{V}_{l}$ for each $l\in[L]$ are defined in (3.2.1).
Proof.
Let $\bm{W}^{*}=(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}_{\bm{\sigma}^{*}}$ be such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})=\|\bm{W}-\bm{W}^{*}\|_{F}$. For ease of exposition, let $\bm{H}_{l}:=\bm{U}_{l-1}^{T}\bm{V}_{l}$ for each $l\in\{2,\dots,L\}$. According to (3.2.1), the $(i,j)$-th block of $\bm{H}_{l}$, denoted by $\bm{H}_{l}^{(i,j)}$, is
By (47) with $\delta_{\sigma}\leq\sigma_{\min}^{*}$ and Lemma 2, we have (36). Substituting (3.2.1) and $\bm{H}_{l}=\bm{U}_{l-1}^{T}\bm{V}_{l}$ into (36) yields
$\left\|\bm{H}_{l}\bm{\Sigma}_{l}^{T}\bm{\Sigma}_{l}-\bm{\Sigma}_{l-1}\bm{\Sigma}_{l-1}^{T}\bm{H}_{l}\right\|_{F}\leq\frac{3\sqrt{2}\sigma^{*}_{\max}}{4\lambda}\|\nabla G(\bm{W})\|_{F},\ \forall l=2,\dots,L.$
Using this and the block diagonal structures of $\bm{\Sigma}_{l}$ in (3.2.1) and $\bm{H}_{l}$ in (49), we obtain
$\ \sum_{j=1}^{p+1}\sum_{i=1,i\neq j}^{p+1}\left\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}-\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\right\|_{F}^{2}$
$\leq$ $\ \sum_{j=1}^{p+1}\sum_{i=1}^{p+1}\left\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}-\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\right\|_{F}^{2}\leq\left(\frac{3\sqrt{2}\sigma^{*}_{\max}}{4\lambda}\right)^{2}\|\nabla G(\bm{W})\|_{F}^{2}.$ (50)
@@
$\left\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}\right\|_{F}\geq\lambda_{\min}^{2}(\bm{\Sigma}_{l}^{(j)})\|\bm{H}_{l}^{(i,j)}\|_{F}\geq\left(\sigma^{*}_{t_{j}}-\frac{\delta_{\sigma}}{3}\right)^{2}\|\bm{H}_{l}^{(i,j)}\|_{F},$ (51)
$\left\|\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\right\|_{F}^{2}\leq\lambda_{\max}\left(\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\right)\|\bm{H}_{l}^{(i,j)}\|_{F}\leq\left(\sigma^{*}_{t_{i}}+\frac{\delta_{\sigma}}{3}\right)^{2}\|\bm{H}_{l}^{(i,j)}\|_{F}.$ (52)
@@
$\quad\left\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}-\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\right\|_{F}\geq\left\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}\right\|_{F}-\left\|\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\right\|_{F}$
[…]
Applying the same argument to the case where $i<j$, we obtain the same result as above. This, togher with (3.2.1), yields the following bound for off-diagonal blocks of $\bm{H}_{l}$:
@@
$\left\|\bm{H}_{l}^{(i,i)}-\bm{P}_{l}^{(i)}\bm{Q}_{l}^{(i)^{T}}\right\|_{F}^{2}$ $=\sum_{k=1}^{g_{i}}\left(1-\sigma_{k}(\bm{H}_{l}^{(i,i)})\right)^{2}\leq\sum_{k=1}^{g_{i}}\left(1-\sigma_{k}^{2}(\bm{H}_{l}^{(i,i)})\right)^{2}$